# Břehová 7, 11519 Prague 1, Czech Republic Francesco Becattini Università di Firenze and INFN Sezione di Firenze, Via G. Sansone 1, I-50019 Sesto Fiorentino (Florence), Italy

### v4

@@ I Introduction
Particularly in very high energy collisions where the chemical potentials are negligible and the hadronization hypersurface $\Sigma$ can be approximated by an isothermal one, the mean spin polarization vector of Dirac fermions such as the $\Lambda$ hyperons with momentum $p$ is given by Becattini et al. (2021b)  [^1] It is important to point out that the formula (1) somewhat differs from others that have been used in other numerical studies Fu et al. (2021); Wu et al. (2022); Jiang et al. (2023a, b); Ribeiro et al. (2024), for a twofold reason. First, the equation (1) is appropriate for an isothermal hadronization hypersurface and not for a general one hence it is specifically applicable to very high energy, where it is a better approximation than the general formula, including temperature gradients (see discussion in Becattini et al. (2021b)). Second, the energy $\varepsilon=p\cdot\hat{t}=p^{0}$ is, in some studies, replaced by $p\cdot u$ where $u$ is the four-velocity field; it appears that this change may lead to significant quantitative differences Chun Shen .:
@@
We adopt the natural units in this work, with $\hbar=c=K=1$. The Minkowskian metric tensor $g$ is ${\rm diag}(1,-1,-1,-1)$; for the Levi-Civita symbol we use the convention $\epsilon^{0123}=1$. We will use the relativistic notation with repeated indices assumed to be saturated.
@@ II Numerical framework
The hydrodynamic evolution is simulated with vHLLE code Karpenko et al. (2014), where the particlization hypersurface is reconstructed with Cornelius subroutine Huovinen and Petersen (2012). This hypersurface is used for the evaluation of $\Lambda$ polarization as well as for conventional Monte-Carlo sampling of hadrons using smash-hadron-sampler code sma (2023). Finally, SMASH Weil et al. (2016) handles rescatterings of the produced hadrons and decays of unstable resonances. We emphasize that spin degrees of freedom are not implemented in SMASH therefore, all the polarization results presented in this work (and so far in the literature) are solely based on formula (1) and the subsequent calculation of polarization transfer in the decay of resonances. The polarization calculations are performed using a dedicated code, hydro-foil, which is publicly available HYDROdynamic Freeze-Out IntegraLs . To conveniently handle all the chain stages, we have created a hybrid model based on the Python programming language and Snakemake Mölder et al. (2021); SMASH-vHLLE hybrid with Python .
@@
As has been mentioned, we have used two Initial State (IS) models in this study. The first one is superMC, numerically implemented from scratch based on the formulae from Shen and Alzhrani (2020); Alzhrani et al. (2022). superMC is based on Glauber geometry with local energy and momentum conservation conditions and provides a $\sqrt{T_{A}T_{B}}$ scaling of the energy density similar to TrENTo Moreland et al. (2015) $p=0$ initial state. The second is a 3D extension of the initial state from Monte Carlo Glauber generator GLISSANDO Rybczynski et al. (2014). With either IS option, we have used the starting time of the fluid stage $\tau_{0}$ and shear viscosity to entropy density ratio $\eta/s$ values that are considered optimal to reproduce basic hadronic observables. In the case of superMC IS, the starting time of the fluid stage is $\tau_{0}=1$ fm/c and, unless otherwise stated, a fixed shear viscosity over entropy density ratio of $\eta/s=0.08$. In the case of 3D GLISSANDO IS, the initial time is $\tau_{0}=0.4$ fm/c, $\eta/s=0.08$ at RHIC energy and $\eta/s=0.1$ at LHC energy, which reflects a somewhat higher temperature range probed in heavy-ion collisions at the LHC. Instead, for the bulk viscosity over entropy density $\zeta/s$, we have used the temperature-dependent parametrization introduced in Schenke et al. (2020), referred to as parametrization III in section V (see eq. 12).
@@
The 3D extension of GLISSANDO initial state is taken from the study in ref.Cimerman et al. (2021), and it follows a parametrization previously reported in Bożek et al. (2015). Differently from ref. Cimerman et al. (2021), we have not computed the normalization of the initial state energy density by identifying the total energy of the fluid with the sum of energies of incoming participants. Instead, we have treated normalization as another free parameter and tuned it to fit charged hadron $dN/d\eta$ distribution. The resulting parameters are reported in Table 2, and figure 2 shows the results of our calculations. We generally have a very good agreement of transverse momentum spectra and pseudo-rapidity distributions of charged hadrons at various centralities but a less good agreement for their elliptic flow coefficient.
@@ IV Polarization at very high energy: results
Simulations have been carried out using averaged initial states, which, as has been discussed in section II, were defined by the models superMC and GLISSANDO. For the superMC IS, the initial state was generated by averaging 20k samples. For GLISSANDO IS, 1k and 3k samples were used. In all cases, we compute polarization of $\Lambda$s in the rapidity window $\eta\in[-1,1]$ to match experimental measurements. Polarization as a function of the azimuthal angle $\langle P_{z}(\phi)\rangle$ and $\langle P_{J}(\phi)\rangle$ are integrated over the $p_{T}$ range $(0,6)$ GeV weighted by the $\Lambda$ spectrum, whereas $\langle P_{z}\sin 2\phi\rangle$ is the mean over the azimuthal angle $\phi$.
@@ V Sensitivity to initial conditions and transport coefficients
We have also studied the dependence of the polarization on the shear viscosity and the bulk viscosity. In figure 7, the transverse and longitudinal components of the spin vector of primary $\Lambda$s with the superMC IS, averaged over the same kinematic intervals as in section II, are shown at RHIC and LHC energies for different constant ratios of $\eta/s$. The transverse component is almost insensitive to the shear viscosity, while the longitudinal component has a limited sensitivity. Interestingly, we notice that while at RHIC energies, a larger viscosity enhances the signal, at LHC, it slightly reduces it.
@@
We have computed the above components for two bulk viscosities: parametrization III of $\zeta/s$ in equation (12) and for vanishing $\zeta/s=0$, see figure 10. In Au-Au collisions at RHIC energy, the contribution from angular velocity is almost vanishing, according to previous findings Karpenko and Becattini (2019) and contrary to the naive lore, which identifies rotation as the main source of polarization. Overall, for parametrization III, the amplitude of both $S_{\sigma}$ and $S_{\theta}$ is smaller than for $\zeta=0$ case, but their sum is approximately zero and the total polarization is almost unaffected. On the other hand, in Pb-Pb collisions at LHC energy, the situation is more complicated. Without bulk viscosity, $S_{\theta}$ seemingly cancels $S_{\sigma}$, and $S_{A_{\omega}}$ dictates the (negative) sign of the polarization harmonic. Turning on the bulk viscosity, not only do $S_{\theta}$ and $S_{\sigma}$ get smaller in magnitude, but also $S_{\omega}$ and $S_{A_{\Xi}}$ significantly change, resulting in the change of sign of the oscillation pattern.

### v5

@@ I Introduction
Particularly in very high energy collisions where the chemical potentials are negligible and the hadronization hypersurface $\Sigma$ can be approximated by an isothermal one, the mean spin polarization vector of Dirac fermions such as the $\Lambda$ hyperons with momentum $p$ is given by Becattini et al. (2021b)  [^1] It is important to point out that the formula (1) somewhat differs from others that have been used in other numerical studies Fu et al. (2021); Wu et al. (2022); Jiang et al. (2023a, b); Ribeiro et al. (2024), for a twofold reason. First, the equation (1) is appropriate for an isothermal hadronization hypersurface and not for a general one hence it is specifically applicable to very high energy, where it is a better approximation than the general formula, including temperature gradients (see discussion in Becattini et al. (2021b)). Second, the energy $\varepsilon=p\cdot\hat{t}=p^{0}$ is, in some studies, replaced by $p\cdot u$ where $u$ is the four-velocity field; this change leads to significant quantitative differences. Indeed, in Ref. Alzhrani et al. (2022) the authors reproduce the correct sign of $P_{z}(\phi)$ without the isothermal decoupling but using $p\cdot u$, and this would not have been the case had they used $p^{0}$ Chun Shen . We stress that the derivation adopted in Ref. Becattini et al. (2021a) requires to use $p^{0}$.:
@@
We adopt the natural units in this work, with $\hbar=c=K=1$. The Minkowskian metric tensor $g$ is ${\rm diag}(1,-1,-1,-1)$; for the Levi-Civita symbol we use the convention $\epsilon^{0123}=1$. We will use the relativistic notation with repeated indices assumed to be summed over.
@@ II Numerical framework
The hydrodynamic evolution is simulated with vHLLE code Karpenko et al. (2014), where the particlization hypersurface is reconstructed with Cornelius subroutine Huovinen and Petersen (2012). This hypersurface is identified as the constant-energy-densty hypersurface $e=0.4$ GeV/fm3, and it is used both for the evaluation of $\Lambda$ polarization as well as for conventional Monte-Carlo sampling of hadrons using smash-hadron-sampler code sma (2023). Finally, SMASH Weil et al. (2016) handles rescatterings of the produced hadrons and decays of unstable resonances. We emphasize that spin degrees of freedom are not implemented in SMASH therefore, all the polarization results presented in this work (and so far in the literature) are solely based on formula (1) and the subsequent calculation of polarization transfer in the decay of resonances. The polarization calculations are performed using a dedicated code, hydro-foil, which is publicly available HYDROdynamic Freeze-Out IntegraLs . To conveniently handle all the chain stages, we have created a hybrid model based on the Python programming language and Snakemake Mölder et al. (2021); SMASH-vHLLE hybrid with Python .
@@
As has been mentioned, we have used two Initial State (IS) models in this study. The first one is superMC, numerically implemented from scratch based on the formulae from Shen and Alzhrani (2020); Alzhrani et al. (2022). superMC is based on Glauber geometry with local energy and momentum conservation conditions and provides a $\sqrt{T_{A}T_{B}}$ scaling of the energy density similar to TrENTo Moreland et al. (2015) $p=0$ initial state. The second is a 3D extension of the initial state from Monte Carlo Glauber generator GLISSANDO Rybczynski et al. (2014), where entropy depositions from participant nucleons are taken to be tilted in space-time rapidity according to Bozek (2012), whereas the depositions from binary scatterings are symmetric in space-time rapidity. With either IS option, we have used the starting time of the fluid stage $\tau_{0}$ and shear viscosity to entropy density ratio $\eta/s$ values that are considered optimal to reproduce basic hadronic observables. In the case of superMC IS, the starting time of the fluid stage is $\tau_{0}=1$ fm/c and, unless otherwise stated, a fixed shear viscosity over entropy density ratio of $\eta/s=0.08$. In the case of 3D GLISSANDO IS, the initial time is $\tau_{0}=0.4$ fm/c, $\eta/s=0.08$ at RHIC energy and $\eta/s=0.1$ at LHC energy, which reflects a somewhat higher temperature range probed in heavy-ion collisions at the LHC. Instead, for the bulk viscosity over entropy density $\zeta/s$, we have used the temperature-dependent parametrization introduced in Schenke et al. (2020), referred to as parametrization III in section V (see eq. 12).
@@
The 3D extension of GLISSANDO initial state is taken from the study in ref.Cimerman et al. (2021), and it follows a parametrization previously reported in Bożek et al. (2015). Differently from ref. Cimerman et al. (2021), we have not computed the normalization of the initial state energy density by identifying the total energy of the fluid with the sum of energies of incoming participants. Instead, we have treated normalization as another free parameter and tuned it to fit charged hadron $dN/d\eta$ distribution. The resulting parameters are reported in Table 2, and figure 2 shows the results of our calculations. We generally have a very good agreement of transverse momentum spectra and pseudo-rapidity distributions of charged hadrons at various centralities but a less good agreement for their elliptic flow coefficient. We attribute the underestimated elliptic flow signal with both initial states to fluid dynamic description with averaged initial state, as compared to event-by-event fluid-dynamic picture in other studies. We also note that at the LHC energy the agreement with the data is better, as the ensuing fluid stage has a longer lifetime, therefore the final flow is somewhat less sensitive to the presence of fluctuations in the initial stage.
@@ IV Polarization at very high energy: results
Simulations have been carried out using averaged initial states, which, as has been discussed in section II, were defined by the models superMC and GLISSANDO. For the superMC IS, the initial state was generated by averaging 20k samples. For GLISSANDO IS, 2k – 15k samples were used, depending on centrality. We conduct the study with event-averaged initial states, to show generic effects of bulk viscosity on polarization observables. We note that, in an event-by-event fluid dynamic description, a fluctuating and spiky initial state will result in a stronger radial flow, which will need to be compensated by a larger bulk viscosity in order for the mean $p_{T}$ to agree with an experimentally measured value. Therefore in a more realistic event-by-event simulation, we expect the effects of bulk viscosity to be even larger than in the present study.
In all cases, we compute polarization of $\Lambda$s in the rapidity window $\eta\in[-1,1]$ to match experimental measurements. Polarization as a function of the azimuthal angle $\langle P_{z}(\phi)\rangle$ and $\langle P_{J}(\phi)\rangle$ are integrated over the $p_{T}$ range $(0,6)$ GeV weighted by the $\Lambda$ spectrum, whereas $\langle P_{z}\sin 2\phi\rangle$ is the mean over the azimuthal angle $\phi$.
@@ V Sensitivity to initial conditions and transport coefficients
We have also studied the dependence of the polarization on the shear viscosity and the bulk viscosity. In figure 7, the transverse and longitudinal components of the spin vector of primary $\Lambda$s with the superMC IS, averaged over the same kinematic intervals as in section IV, are shown at RHIC and LHC energies for different constant ratios of $\eta/s$ whereas the bulk viscosity is given by the parametrization III as explained in sec. II. The transverse component is almost insensitive to the shear viscosity, while the longitudinal component has a limited sensitivity. Interestingly, we notice that while at RHIC energies, a larger viscosity enhances the signal, at LHC, it slightly reduces it.
@@
We have computed the above components for two bulk viscosities: parametrization III of $\zeta/s$ in equation (12) and for vanishing $\zeta/s=0$, see figure 10. In Au-Au collisions at RHIC energy, the contribution from angular velocity is almost vanishing, according to previous findings Karpenko and Becattini (2019) and contrary to the naive lore, which identifies rotation as the main source of polarization. Overall, for parametrization III, the amplitude of both $S_{\sigma}$ and $S_{\theta}$ is smaller than for $\zeta=0$ case, but their sum is approximately zero and the total polarization is almost unaffected. On the other hand, in Pb-Pb collisions at LHC energy, the situation is more complicated. Without bulk viscosity, $S_{\theta}$ seemingly cancels $S_{\sigma}$, and $S_{A_{\Xi}}$ dictates the (negative) sign of the polarization harmonic. Turning on the bulk viscosity, not only do $S_{\theta}$ and $S_{\sigma}$ get smaller in magnitude, but also $S_{\omega}$ and $S_{A_{\Xi}}$ significantly change, resulting in the change of sign of the oscillation pattern.